How many tiles whose length and breadth are 12 cm and 5 cm respectively will be needed to fit in a rectangular region whose length and breadth are respectively?
100 cm and 144 cm

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Area of rectangle = 100× 144
= 14400 cm
Area of one tile = 5× 12
= 60 cm^2
Number of tiles = (Area of rectangle) / (Area of one tile)
= 14400 / 60
= 240
Hence, 240 tiles are needed

Thanks